The contentDM server was renamed from digitalassets to digitalcollections. This change was made so the name of the server better reflects the services on the server.
Automatic startup scripts were added for memcached, sqlrelay, ezproxy, and mysql on the University’s web server, www.ups.edu
Because the web server’s disk filled up again as on March 2, the we server was restarted this morning.
ELI Internet Customer Support Notice:
Electric Lightwave, Inc will be installing a new
PA-2T3 card in the router; GW01:STTL in Seattle, WA. We will be doing this on March 26, 2004 at 5:00am PST. During this maintenance, customers may experience an outage of approximately 30 minutes. We apologize for any inconvenience this may cause.
Start: The outage will occur on March 26, 2004 at 5:00am PST.
Duration: The outage will last approximately 30 minutes.
An unknown malfuction in multicast routing caused a network slowdown (and failure in some areas of campus). Multicast routing was disabled to solve the problem.
For background on this problem, see the January 12 weblog entry.
A new domain, mail.ups.edu, was defined with an MX record pointing mail.ups.edu to the antivirus gateway (AVG). This allows us to deliver messages destined for addresses of the form firstname.lastname@example.org. We will leave this in place until 1 July.
Beginning on Monday March 22, 2004 we will begin upgrading the University mailer daemons. The daemons are the applications that allow users to send, receive and manage their e-mail. There will be brief interruptions during the week of March 22, 2004.
If you encounter errors or problems, users are advised to wait ten minutes and try again. If the problem is not corrected, please contact the OIS Help Desk, email@example.com or X8585.
I checked the C-BORD Oyssey backup process this moring to inspect my changes. It was determined that the modifications made on 3/18 were not executed. Upon further inspection it was determined that the original file did not use the correct command-line version switches for the backup log. This was modified and we wait for the scheduled process to run again.
This morning there were questions posed by the operators about the status of the ntbackup process on the C-BORD Odyssey server around the location of the backup log file.
During the course of my investigation, I discovered that the use of the %Odyssey% variable does not work in all situations. So I modified the tapebackup.cmd file to use the absolute path contained in the %Odyssey% variable.
Upgraded IOS on 6509 and 3640 on 3/15/2004. 6509 upgraded from IP W/SSH/3DES12.2(14)SX1 to 12.2(17d).SXB. 3540 upgraded from IP 12.2(T1) to 12.23.
Upgraded 3005 vpn to ver 4.1.2 on 3/10/2004.
The University mail server began to timeout clients and slow down at approximately noon today. The cause of the problem is unclear.
The number of process running on the server appearred to be at a normal load average. The servers response from secure shell was slow. It appearred as though the system was having a problem allocating resources for the processes running, but memory and disk space were both available.
We shut down process to try and identify the cause of the slowness. Imapd and Ipop3d were disabled with no luck. The web server was shutdown, no luck. Then mailman and sendmail. At this point things improved.
I disabled the webmail interface and brought sendmail and mailman backup. I then brought the web server backup. The server still appeared to be responding in a slower than normal, but somewhat timely manner. I restarted the ipop3d daemon and about five minutes later the imapd daemon. The server still looked good–slower than normal, but somewhat timely. After restarting the webmail interface things when into the tank.
We tried restarting the webmail server between 3:00 and 3:30 pm, but the unbareable slowness remained.
We then rebooted the mail server at about 3:40pm.
The server was back to normal after the reboot.
At 10:30 am load testing of the media server was performed. This testing involved an audio file that was over 19 minutes long (19:37) and 69 workstations with the real media client.
The clients were located in Wyatt 203, McIntyre 324, and basement of the library with a few clients in the I-Commons. The majority of these machines were connected to campus network via cabletron 6000 switches.
The test lasted for the length of the audio file. During that time, no adverse network activity was observed. The server was relatively uneffected as well. At most, 18% of the cpu was consumed for a brief period of time. No service related calls were logged by the Help Desk concerning network issues during the testing. It should be noted that this test was performed during Spring Break.
A new license key was generated, and was put into place when the web server was restarted today. Wu-ftpd was stopped.
The disk on the University’s web server filled up sometime last week, causing several cron processes to hang and “zombify”. In order to clear up these problems, the web server was rebooted.
The license key for the SurgeFTP server was apparently incorrectly generated. As a consequence, the server was running on the demo key, which expired today. SurgeFTP was stopped, and WU-ftpd was restarted so that people can upload their files. We are writing